Shift Summary, DAY Nov 15, 1999

Herman B. White (hwhite@fnal.gov)
Mon, 15 Nov 1999 15:58:14 -0600 (CST)

Shift Summary for DAY Shift, 15/Nov/99

On Shift: S.E. Turner, HBW

Hours of beam available at all intensities: 8.4 hrs
Hours of beam available at nominal intensity: 8.0 hrs
Hours of E799 data (special runs not included): 6.5 hrs
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time:

tape changes, trigger errors, CSI low-voltage power failure.

========================================================================
Runs ended this shift:
Run 14975 Beam_799 tapes full
Run 14978 BEAM_799, some problems only a few spills
Run 14979 BEAM_799, OK up to spill 252,

Tapes brought to FCC this shift: none

========================================================================
Accesses during shift: one
Investigating low voltage sector-9 failure

========================================================================
Overview of shift:

A few tape problems with identifying split tapes and starting of
the initial run. Most of the shift was smooth, and no real problems until
the end of the shift with a failed CsI sector nine power supply. A
solution was possibly found for the 2Hz oscillation problem. Apparently an
identified bit set in the accelerator and associated with the QXR can
cause the problem.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

14975 151-153 07:50 Trigger error/ktevana errors ordinal number
mismatch in 2X bananas. 3 kfrends for fix

14975 08:10 Run Ends

14978 08:50 Begin Run after reloading KQ and reset Bananas
get L3 accept fraction warning, find ktev2 dead
reset 167 and fresh start (overwrite 14978 1-spill)
14979 09:10 Begin Run after reloading KQ
09:36 Replaced split upm-029 after cleaning, old tape
kept and labled "07:00 -9:30AM"
14979 53 10:14 MCR calls and resets QXR as part of study of
2Hz problem. They believe problem is solved.
77 11:38 problem and solution identified for 2Hz problem
252 14:44 SDAQ shows error on CSI-Sector-9 Low voltage supply
14:55 Access made, burning ordor observed, supply not
reset and JAWS and Sydney T. investigating
15:03 Run ended

========================================================================
Follow up Notes on Previous Shift
Banana problems investigatied and YTF problem stable for 8hrs.


========================================================================
Items to Follow up on Next Shift